can you help me solve this?
4(2w+5)=12w-9

Answer by checkley77(12844)

4(2w+5)=12w-9
8W+20=12W-9
8W-12W=-9-20
-4W=-29
W=-29/-4
W=29/4 ANSWER.
PROOF
4(2*29/4+5)=12*29/4-9
4(29/2+5)=3*29-9
4(29+10)/2=87-9
2(39)=78
78=78